Show that the functions 1, cos2 x, sin2 x are linearly dependent .


1
Expert's answer
2022-04-11T16:38:12-0400

Assume that there are complex numbers "a_1,a_2,a_3" not all equal to zero satisfying: "a_1+a_2cos\,2x+a_3sin\,2x=0". Choose different values of "x:" "x=0,x=\frac{\pi}{4},x=\frac{\pi}{3}". We get the following equations: "a_1+a_2=0,\,\, a_1+a_3=0=0\,\,a_1-\frac12a_2+a_3\frac{\sqrt{3}}{2}=0". From first and second equations we get: "a_2=a_3=-a_1" . Third equation implies: "a_3=0". Thus, all numbers are zero. We came to contradiction and received that all functions are linearly independent.